OpenStudy (anonymous):

PLEASE HELP!!! I WILL GIVE A MEDAL!! Find the sum of the given infinite geometric series. A1=32,r=-0.06 A.32.6 B.34.04 C.30.19 D.48.67 A1=-8,r=-0.02 A.-7.84 B.-8.16 C.-7.21 D.-5.38 A1=22,r=-0.07 A.36.29 B.20.56 C.23.66 D.22.7

OpenStudy (here_to_help15):

a_1/(1-r) so 32/(1-(-.06)) = 32/(1+.06) = 32/1.06